40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Nov 7 14:39:49 2013, Jenne, Update, LSC, Lock Acquisition Game Plan LockAcquisition_7Nov2013.pdf
    Reply  Fri Nov 8 18:12:21 2013, Jenne, Update, LSC, PRFPMI: Not crossing any resonances 2f_resonances_Xarm_CARMsweep.pngAllModFreq_resonances_Xarm_CARMsweep.png
    Reply  Tue Nov 12 16:49:22 2013, Jenne, Update, LSC, Xend QPD and Whitening board pulled IMG_1262.JPGIMG_1260.JPGIMG_1264.JPGIMG_1276.JPG
       Reply  Wed Nov 20 17:05:15 2013, Jenne, Update, LSC, Xend QPD and Whitening board replaced 
       Reply  Wed Dec 18 00:26:15 2013, Jenne, Update, LSC, Xend QPD schematic investigation 
          Reply  Wed Dec 18 11:32:34 2013, rana, Update, LSC, Xend QPD schematic investigation 
    Reply  Wed Nov 13 01:35:40 2013, Jenne, Update, LSC, PRM motion causing trouble? xarmRIN_vs_CARM.pngASportOverlap_moveGaussianBeams.png
       Reply  Wed Nov 13 09:31:15 2013, Gabriele, Update, LSC, PRM motion causing trouble? 
          Reply  Wed Nov 13 19:22:58 2013, Jenne, Update, LSC, PRM motion correlated to intracavity power POPDC_sensors_30oct2013.pngAOMmodulated_POPDC_OSEM.pdf
             Reply  Thu Nov 14 02:50:43 2013, Jenne, Update, LSC, PRM oplev measured and modeled TF PRM_OL_TF_model.pngPRM_OLPIT_TF.pdfPRM_OLYAW_TF.pdf
                Reply  Mon Nov 18 21:02:54 2013, Jenne, Update, LSC, PRM oplev measured and modeled TF PRM_OLG_NewOld.pngPRM_oplev_improvement.pdf
                   Reply  Mon Nov 18 21:40:24 2013, Koji, Update, LSC, PRM oplev measured and modeled TF 
                      Reply  Tue Nov 19 14:47:44 2013, Jenne, Update, LSC, PRM oplev measured and modeled TF 
             Reply  Thu Nov 14 02:55:26 2013, rana, Update, SUS, PRM motion correlated to intracavity power 
                Reply  Thu Nov 14 11:41:19 2013, Steve, Update, SUS, PRM sensors effected by IR IRon_off.png
                   Reply  Thu Nov 14 14:35:12 2013, Steve, Update, SUS, IR effect on MC and PRM sensors MC_PRM_IReffect.png
                      Reply  Fri Dec 6 10:03:07 2013, Steve, Update, SUS, IR effect on MC sensors only 10minPMCnotLocked.png6Dec2013.png
Message ID: 9401     Entry time: Mon Nov 18 21:02:54 2013     In reply to: 9382     Reply to this: 9404
Author: Jenne 
Type: Update 
Category: LSC 
Subject: PRM oplev measured and modeled TF 

I have created a new filter for the PRM oplev damping loops.  The biggest change is an increase in the gain between 0.4 - 7 Hz.

Here is a plot of the old, and my new modelled open loop gain:

PRM_OLG_NewOld.png

When I look at my step and impulse response time series, the notches for the bounce and roll were causing some ringing, so for now they are turned off, both in the model and in the real time system.  Also, the "OLG orig" trace has a 4th order elliptic lowpass at 75 Hz, but the real system had a 4th order elliptic low pass at 35 Hz.  When we use 35 Hz in the model, we get lots of ringing.  So, we have moved both model and real system to 55 Hz 4th order elliptic low passes.  Also, also, we haven't been using the 3.3 Hz resonant gain, so I removed that from the modelled loop.

I have put the "boost" for the .4-7 Hz emphasis into FM 7 of the PRM oplev filters.  I also removed several old filters that are never used.  So, for now, the PRM oplevs should have engaged:  FM 1, 7, 9. Pitch gain is +5, yaw gain is -9.  We can consider re-implementing the bounce-roll notches, and the stack resgain if it looks like those are getting rung up, and causing trouble.

Here is a set of spectra, showing the improvement.  It's unclear why yaw is worse than pitch below 4Hz, and why pitch is so much worse than yaw between 4-15 Hz, however for each of pitch and yaw, the before (reference pink and cyan traces) is higher than the improved (dark red, dark blue traces) between a few tenths of a Hz up to 3ish Hz.  And, we're not causing more noise elsewhere.  We do want to monitor to make sure we're not ringing up the bounce and roll modes, but for now they seem fine.

PRM_oplev_improvement.pdf

ELOG V3.1.3-